OpenStudy (explorerhills):

Why is the range of theta, (in dot product of two vectors) 0 to pi ?

OpenStudy (explorerhills):

@ganeshie8 @parthkohli @Mehek14

OpenStudy (faiqraees):

Theta is the angle between the two vectors and it can be any angle value

OpenStudy (faiqraees):

so 0 to 2 pi

OpenStudy (welshfella):

I think you would tend to quote the obtuse angle though, rather than the reflex angle. So in that respect the range is 0 to pi.

OpenStudy (error1603):

http://openstudy.com/users/ganeshie8

OpenStudy (error1603):

@parthkohli

OpenStudy (error1603):

@Mehek14

OpenStudy (explorerhills):

@ganeshie8

ganeshie8 (ganeshie8):

\[\theta = \arccos\left(\dfrac{v\cdot w}{|v||w|}\right)\] what is the range of \(\arccos(x)\) ?

OpenStudy (explorerhills):

0 to pi..

OpenStudy (explorerhills):

okay can we say that we can draw a plane given any two vectors , hence the maximum angle between then can be pi ?

ganeshie8 (ganeshie8):

how are they related ?

OpenStudy (explorerhills):

We use dot product to calculate the angle between two vectors, so i thought about putting it this way..?